OpenStudy (anonymous):

Solve -11-_x+30)=-41-2(5x-18)

OpenStudy (anonymous):

CALCULATOR NEEDED metacalc.com

OpenStudy (anonymous):

Is there really an underline?

OpenStudy (anonymous):

no it suposed to be (x+30) sorry

OpenStudy (anonymous):

oh that makes sense now

OpenStudy (anonymous):

x= 4

OpenStudy (anonymous):

-11-x-30=-41-10x+36 -41-x = -5-10x 9x = 36 x=4

OpenStudy (anonymous):

thanks